OpenStudy (anonymous):

Show an equation and a solution for the problem. How much water must be added to 12 L (liters) of a 40% solution of alcohol to obtain a 30% solution?

OpenStudy (cwrw238):

let x be the total volume after adding the water then as the amount of alcohol stays constant 12 * 0.4 = x * 0.3 x = 4.8 / 0.3 = 16 litres so amount added = what?

OpenStudy (jackellyn):

The amount of alcohol in the solution is 40% of 12L, or 0.40*12L=4.8L If that amount is to be 30% of the expected volume (in liters) of 12+x, then 0.30*(12+x)=4.8 3.6+0.30x=4.8 --> 0.30x=1.2 --> x=1.2/0.30 --> x=4.0 You need to add 4.0L water.

OpenStudy (cwrw238):

volume was 12 now its 16

OpenStudy (cwrw238):

- as jackelyn said

OpenStudy (anonymous):

thanksss

OpenStudy (cwrw238):

yw